# Taehoon Yoon     Yunhong Min     Kyeongmin Yeo     Minhyuk Sung KAIST

### v2

@@ 1 Introduction
A straightforward baseline for generating initial particles is the Top-$K$-of-$N$ strategy: drawing multiple samples from the standard Gaussian prior and selecting those with the highest reward scores. Though effective, this naive approach offers limited improvement in subsequent SMC due to its reliance on brute-force sampling. Motivated by these limitations, we explore Markov Chain Monte Carlo (MCMC) [30, 31, 32, 33, 34, 35] methods based on Langevin dynamics, which are particularly well-suited to our setting since we sample from the initial posterior distribution, whose form is known. Nevertheless, applying MCMC in our problem presents unique challenges: the exploration space is extremely high-dimensional ($e.g.$, 65,536 for FLUX [28]), posing significant challenges for conventional MCMC methods. In particular, the Metropolis–Hastings (MH) accept-reject mechanism, when used with standard Langevin-based samplers, becomes ineffective in such high-dimensional regimes, as the acceptance probability rapidly diminishes and most proposals are rejected.
@@
In our experiments, we evaluate three reward alignment tasks: layout-to-image generation (placing objects in designated bounding boxes within the image), quantity-aware generation (aligning the number of objects in the image with the specified count), and aesthetic-preference generation (enhancing visual appeal). We compare our $\Psi$-Sampler against the base SMC method [14] with random initial particle sampling, SMC combined with initial particle sampling via Top-$K$-of-$N$, ULA, and MALA, as well as single-particle methods [39, 40]. Across all tasks, $\Psi$-Sampler consistently achieves the best performance in terms of the given reward and generalizes well to the held-out reward, matching or surpassing existing baselines. Its improvement over the base SMC method highlights the importance of posterior-based initialization, while its outperformance over ULA and MALA further confirms the limitations of these methods in extremely high-dimensional spaces.
@@ 3.2 Inference-Time Reward Alignment Using Score-Based Generative Models
where $\mathbf{x}_{0|t}:=\mathbb{E}_{\mathbf{x}_{0}\sim p_{0|t}}[\mathbf{x}_{0}]$ denotes Tweedie’s formula [18], representing the conditional expectation under $p_{0|t}:=p(\mathbf{x}_{0}|\mathbf{x}_{t})$. Under the same approximation, the transition kernel satisfying the optimality condition is given by:
@@ 4 $\Psi$-Sampler: pCNL-Based Initial Particle Sampling
In this work, we propose $\Psi$-Sampler, a framework that combines efficient initial particle samping with SMC-based inference-time reward alignment for score-based generative models. The initial particles are sampled using the Preconditioned Crank–Nicolson Langevin (pCNL) algorithm, hence the name PSI ($\Psi$)PCNL-based Sampling for Inference-time scaling. The key idea is to allocate computational effort to the initial particle selection by sampling directly from the posterior distribution defined in Eq. 4. This reward-informed initialization ensures that the particle set is better aligned with the target distribution from the outset, resulting in improved sampling efficiency and estimation accuracy in the subsequent SMC process.
While the unnormalized density of the posterior distribution in Eq. 4 has an analytical form, drawing exact samples from it remains challenging. A practical workaround is to approximate posterior sampling via a Top-$K$-of-$N$ strategy: generate $N$ samples from the prior, and retain the top $K$ highest-scoring samples as initial particles. This variant of Best-of-$N$ [64, 65, 66] resembles rejection sampling and serves as a crude approximation to posterior sampling [67, 68]. We find that even this simple selection-based approximation leads to meaningful improvements. But considering that sampling space is high-dimensional, one can adopt Markov Chain Monte Carlo (MCMC) [30, 31, 32, 33, 34, 35] which is known to be effective at sampling from high-dimensional space. In what follows, we briefly introduce Langevin-based MCMC algorithms that we adopt for posterior sampling.
@@ 4.1 Background: Langevin-Based Markov Chain Monte Carlo Methods
Langevin-based MCMC refers to a class of samplers that generate proposals by discretizing the Langevin dynamics, represented as stochastic differential equation (SDE),
@@ 4.3 Initial Particle Sampling
In Fig. 1, we present a 2D toy experiment comparing SMC performance when initializing particles from the prior versus the posterior. We train a simple few-step score-based generative model on a synthetic dataset where the clean data distribution $p_{0}$ is a 6-mode Gaussian Mixture Model (GMM), shown as red dots in Fig.1 (A). The prior distribution is shown in blue, and the gray lines depict sampling trajectories during generation. We define a reward function that assigns high scores to samples from only a subset of the GMM modes, yielding a target distribution at $t=0$ (Eq. 3), as illustrated in Fig. 1 (B) (red dots). The corresponding optimal initial distribution—the posterior at $t=1$ (Eq. 4)—is shown as blue dots in Fig. 1 (B). We compare (C) standard SMC with prior sampled particles, (D) SMC with posterior samples from MALA, and (E) our $\Psi$-Sampler. All settings use the same total number of function evaluations (NFE). Prior-based SMC (C) uses 100 NFE; MALA+SMC and $\Psi$-Sampler allocate 50 NFE for MCMC and use fewer particles for SMC. While MALA-based initialization method (D) significantly improves alignment with the target distribution (red dots in (A)) over prior-based method (C), some modes remain underrepresented. In contrast, $\Psi$-Sampler (E) provides tighter alignment with the target distribution and better posterior coverage, even in this simple 2D setting, illustrating its effectiveness in sampling high-quality samples.
@@ 5.1 Experiment Setup
We validate our approach across three applications: layout-to-image generation, quantity-aware generation, and aesthetic-preference image generation. In our experiments, the held-out reward refers to an evaluation metric that is not accessible during generation and is used solely to assess the generalization of the method. Full details on reward models and experimental setups are provided in the Appendix D.
@@
  • SMC & Initial Particles from Posterior: We evaluate four posterior-based initialization strategies: Top-$K$-of-$N$, ULA, MALA, and $\Psi$-Sampler. ULA and MALA use a small step size (0.05) to ensure non-zero acceptance, while $\Psi$-Sampler employs a larger step size (0.5) for improved performance.
We use 25 denoising steps for SMC-based methods and 50 for single-particle methods to compensate their limited exploration. For SMC-based methods, we match the total number of function evaluations (NFE) across all methods, allocating part of the budget to initial particle sampling for posterior-based cases. Prior-based methods use 40 particles, while posterior-based methods use 20, allocating half of the NFE to initialization. We use 1,000 NFE for layout-to-image generation and quantity-aware image generation, and 500 for the aesthetic-preference image generation. For the aesthetic-preference image generation, we use half the number of particles in both settings to reflect the halved NFE. We used FLUX [28] as the pretrained score-based generative model. Full experimental details are provided in the Appendix D.
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-$N$ ULA MALA $\Psi$-Sampler
@@
[…]
@@ 5.2 Quantitative Results
We present quantitative results in Tab.1. Across all tasks, $\Psi$-Sampler consistently achieves the best performance on the given reward and strong generalization to held-out rewards. For SMC-based methods, sampling particles from the posterior distribution yields significant improvements over those that sample directly from the prior, highlighting the importance of posterior-informed initialization. This improvement is particularly notable in complex tasks where high-reward outputs are rare, such as layout-to-image generation and quantity-aware generation. For example, in quantity-aware generation, negative smooth L1 loss improves from 1.804 with TDS (base SMC) to 1.077 with Top-$K$-of-$N$ and further to 0.850 with our $\Psi$-Sampler. Similarly, for layout-to-image generation, mIoU increases from 0.417 (TDS) to 0.425 with Top-$K$-of-$N$ and 0.467 with $\Psi$-Sampler.
@@
In contrast, initializing with ULA or MALA yields only marginal gains or even degraded performance, due to the lack of Metropolis-Hastings correction in ULA and the limited exploration capacity of MALA in high-dimensional spaces. Single-particle methods consistently underperform compared to SMC-based methods.
Ablation Study.
We conduct an ablation study that examines how performance varies under different allocations of a fixed total NFE between the initial particle sampling stage (via Top-$K$-of-$K$ or MCMC) and the subsequent SMC stage; full results and analysis are provided in Appendix F.
We additionally provide quantitative and qualitative results on all three applications using another score-based generative model, SANA-Sprint [29] in Appendix G.
@@ 5.3 Qualitative Results
From the first and second rows of Fig. 2, we observe that baseline methods often fail to place objects correctly within the specified bounding boxes or generate them in entirely wrong locations. For instance, in the first row, most baselines fail to position the bird accurately, and in the second row, none correctly place the car. For quantity-aware generation, the fourth row shows the counted results corresponding to the third row. While $\Psi$-Sampler successfully generates the target number of blueberries in an image, the baselines exhibit large errors—Top-$K$-of-$N$ comes closest but still misses some. In rows 5 and 6, only $\Psi$-Sampler correctly generates the target number of coins. In the aesthetic preference task, although all methods produce realistic images, $\Psi$-Sampler generates the most visually appealing image with the highest aesthetic score. Additional qualitative examples are provided in the Appendix H.
@@ 5.4 Evaluation of Initial Particles
In Fig.4, we compare MALA and pCNL on the layout-to-image generation task across varying step sizes using four metrics: acceptance probability, reward (mIoU via GroundingDINO [72]), held-out reward (Salience DETR [74]), and sample diversity (LPIPS MPD [79]). All metrics are directly computed from the Tweedie estimates [18] of MCMC samples, before the SMC stage. As the step size increases, MALA’s acceptance probability rapidly drops to near-zero, while pCNL maintains stable acceptance probability. Larger step sizes generally improve reward scores, with performance tapering off at excessively large steps. Held-out reward trends mirror this pattern, suggesting that the improvements stem from genuinely higher-quality samples rather than reward overfitting [60, 20]. Although LPIPS MPD slightly declines with increasing step size due to reduced acceptance, pCNL at step size 2.0 maintains diversity on par with MALA at 0.05. Additional results for other tasks are included in the Appendix E.
@@ Appendix A Reward Alignment with Stochastic Optimal Control
which can be proved [20, 52] using Girsanov theorem and martingale property $\hat{\text{o}}$ integral.
@@ Appendix C Acceptance Probability of MALA and pCNL
[…]
@@
[…]
@@
[…]
@@ Appendix D Experimental Setup and Details
We use FLUX-Schnell [28] as the score-based generative model for our method and all baselines. Although FLUX is a flow-based model, SMC-based inference-time reward alignment can be applied by reformulating the generative ODE as an SDE [7, 13], as described in Sec. 3.1. Further to ensure diversity between samples during SMC, we applied Variance Preserving (VP) interpolant conversion [13, 54]. Apart from FLUX, we additionally report quantitative and qualitative results using another score-based generative model, SANA-Sprint [29], in Appendix G. These results demonstrate that the our claims are not tied to a specific model architecture, and additionally highlighting the robustness of $\Psi$-Sampler.
We use 25 denoising steps for all SMC-based methods and 50 for single-particle methods to compensate for their reduced exploration capacity. To ensure fair comparison, we fix the total number of function evaluations (NFE) across all SMC variants—1,000 for layout-to-image and quantity-aware generation tasks, and 500 for aesthetic-preference image generation. For aesthetic-preference image generation, we used half the NFE compared to other tasks because we found it sufficient for performance convergence. For methods that sample from the posterior, half of the NFE is allocated to the initial sampling stage, resulting in 20 particles during SMC, while prior-based methods use all NFE for SMC with 40 particles. For the aesthetic-preference image generation, we use half the number of particles in both settings to reflect the halved NFE. The ablation study comparing the performance on varying NFE allocation is provided in Appendix F. For all experiments, we used NVIDIA A6000 GPUs with 48GB VRAM.
Appendix E Additional Evaluation Results of MALA and pCNL Initializations under Varying Step Size

### v3

@@ 1 Introduction
A straightforward baseline for generating initial particles is the Top-$K$-of-NN strategy: drawing multiple samples from the standard Gaussian prior and selecting those with the highest reward scores. Though effective, this naive approach offers limited improvement in subsequent SMC due to its reliance on brute-force sampling. Motivated by these limitations, we explore Markov Chain Monte Carlo (MCMC) [30, 31, 32, 33, 34, 35] methods based on Langevin dynamics, which are particularly well-suited to our setting since we sample from the initial posterior distribution, whose form is known. Nevertheless, applying MCMC in our problem presents unique challenges: the exploration space is extremely high-dimensional ($e.g.$, 65,536 for FLUX [28]), posing significant challenges for conventional MCMC methods. In particular, the Metropolis–Hastings (MH) accept-reject mechanism, when used with standard Langevin-based samplers, becomes ineffective in such high-dimensional regimes, as the acceptance probability rapidly diminishes and most proposals are rejected.
@@
In our experiments, we evaluate three reward alignment tasks: layout-to-image generation (placing objects in designated bounding boxes within the image), quantity-aware generation (aligning the number of objects in the image with the specified count), and aesthetic-preference generation (enhancing visual appeal). We compare our $\Psi$-Sampler against the base SMC method [14] with random initial particle sampling, SMC combined with initial particle sampling via Top-$K$-of-NN, ULA, and MALA, as well as single-particle methods [39, 40]. Across all tasks, $\Psi$-Sampler consistently achieves the best performance in terms of the given reward and generalizes well to the held-out reward, matching or surpassing existing baselines. Its improvement over the base SMC method highlights the importance of posterior-based initialization, while its outperformance over ULA and MALA further confirms the limitations of these methods in extremely high-dimensional spaces.
@@ 3.2 Inference-Time Reward Alignment Using Score-Based Generative Models
where $\mathbf{x}_{0|t}:=\mathbb{E}_{\mathbf{x}_{0}\sim p_{0|t}}[\mathbf{x}_{0}]$ denotes Tweedie’s formula [18], representing the conditional expectation under $p_{0|t}:=p(\mathbf{x}_{0}|\mathbf{x}_{t})$. Under the same approximation, the transition kernel satisfying the optimality condition is approximated by:
@@ 4 $\Psi$-Sampler: pCNL-Based Initial Particle Sampling
In this work, we propose $\Psi$-Sampler, a framework that combines efficient initial particle samping with SMC-based inference-time reward alignment for score-based generative models. The initial particles are sampled using the Preconditioned Crank–Nicolson Langevin (pCNL) algorithm, hence the name PCNL-based initial particle sampling followed by SMC-based Inference-time reward alignment. The key idea is to allocate computational effort to the initial particle selection by sampling directly from the posterior distribution defined in Eq. 4. This reward-informed initialization ensures that the particle set is better aligned with the target distribution from the outset, resulting in improved sampling efficiency and estimation accuracy in the subsequent SMC process.
While the unnormalized density of the posterior distribution in Eq. 4 has an analytical form, drawing exact samples from it remains challenging. A practical workaround is to approximate posterior sampling via a Top-$K$-of-NN strategy: generate $N$ samples from the prior, and retain the top $K$ highest-scoring samples as initial particles. This variant of Best-of-$N$ [64, 65, 66] resembles rejection sampling and serves as a crude approximation to posterior sampling [67, 68]. We find that even this simple selection-based approximation leads to meaningful improvements. But considering that sampling space is high-dimensional, one can adopt Markov Chain Monte Carlo (MCMC) [30, 31, 32, 33, 34, 35] which is known to be effective at sampling from high-dimensional space. In what follows, we briefly introduce Langevin-based MCMC algorithms that we adopt for posterior sampling.
@@ 4.1 Background: Langevin-Based Markov Chain Monte Carlo Methods
Langevin‑based MCMC refers to a class of samplers that generate proposals by discretizing the Langevin dynamics, represented as stochastic differential equation (SDE),
@@ 4.3 Initial Particle Sampling
In Fig. 1, we present a 2D toy experiment comparing SMC performance when initializing particles from the prior versus the posterior. We train a simple few-step score-based generative model on a synthetic dataset where the clean data distribution $p_{0}$ is a 6-mode Gaussian Mixture Model (GMM), shown as red dots in Fig.1 (A). The prior distribution is shown in blue, and the gray lines depict sampling trajectories during generation. We define a reward function that assigns high scores to samples from only a subset of the GMM modes, yielding a target distribution at $t=0$ (Eq. 3), as illustrated in Fig. 1 (B) (red dots). The corresponding optimal initial distribution—the posterior at $t=1$ (Eq. 4)—is shown as blue dots in Fig. 1 (B). We compare (C) standard SMC with prior sampled particles, (D) SMC with posterior samples from MALA, and (E) our $\Psi$-Sampler. All settings use the same total number of function evaluations (NFE). Prior-based SMC (C) uses 100 NFE; MALA+SMC and $\Psi$-Sampler allocate 50 NFE for MCMC and use fewer particles for SMC. While MALA-based initialization method (D) significantly improves alignment with the target distribution (red dots in (B)) over prior-based method (C), some modes remain underrepresented. In contrast, $\Psi$-Sampler (E) provides tighter alignment with the target distribution and the posterior distribution, illustrating its effectiveness in sampling high-quality samples. Full experimental details are provided in Appendix E.
@@ 5.1 Experiment Setup
We validate our approach across three applications: layout-to-image generation, quantity-aware generation, and aesthetic-preference image generation. In our experiments, the held-out reward refers to an evaluation metric that is not accessible during generation and is used solely to assess the generalization of the method. Full details for each application are provided in Appendix D.
@@
  • SMC & Initial Particles from Posterior: We evaluate four posterior-based initialization strategies: Top-$K$-of-NN, ULA, MALA, and $\Psi$-Sampler. ULA and MALA use a small step size (0.05) to ensure non-zero acceptance, while $\Psi$-Sampler employs a larger step size (0.5) for improved performance (See Sec. 5.4).
We use 25 denoising steps for SMC-based methods and 50 for single-particle methods to compensate for their limited exploration. For SMC-based methods, we match the total number of function evaluations (NFE) across all methods, allocating half of the budget to initial particle sampling for posterior-based methods. We use FLUX [28] as the pretrained score-based generative model. Full experimental details are provided in Appendix D.
FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler
@@
Single Particle SMC-Based Methods Tasks Metrics Sampling from Prior Sampling from Posterior DPS [39] FreeDoM [40] TDS [14] DAS [12] Top-$K$-of-NN ULA MALA $\Psi$-Sampler Layout to Image $\text{GroundingDINO}$$^{\dagger}$ [72] $\uparrow$ 0.166 0.177 0.417 0.363 0.425 0.370 0.401 0.467 mIoU [74] $\uparrow$ 0.215 0.229 0.402 0.342 0.427 0.374 0.401 0.471 ImageReward [77] $\uparrow$ 0.705 0.713 0.962 0.938 0.957 0.838 0.965 1.035 VQA [78] $\uparrow$ 0.684 0.650 0.794 0.784 0.855 0.783 0.789 0.810 Quantity Aware $\text{T2I-Count}$$^{\dagger}$ [73] $\downarrow$ 14.187 15.214 1.804 1.151 1.077 3.035 1.601 0.850 MAE [75] $\downarrow$ 15.7 15.675 5.3 4.175 3.675 4.825 3.575 2.925 Acc (%) [75] $\uparrow$ 0.0 0.0 27.5 15.0 12.5 22.5 25.0 32.5 ImageReward [77] $\uparrow$ 0.746 0.665 0.656 0.507 0.752 0.743 0.742 0.796 VQA [78] $\uparrow$ 0.957 0.953 0.943 0.907 0.960 0.943 0.941 0.951 Aesthetic Preference $\text{Aesthetic}$$^{\dagger}$ [76] $\uparrow$ 6.139 6.310 6.853 6.935 6.879 6.869 6.909 7.012 ImageReward [77] $\uparrow$ 1.116 1.132 1.135 1.166 1.133 1.100 1.155 1.171 VQA [78] $\uparrow$ 0.968 0.959 0.970 0.970 0.961 0.961 0.952 0.963
@@ 5.2 Quantitative Results
We present quantitative results in Tab.1. Across all tasks, $\Psi$-Sampler consistently achieves the best performance on the given reward and strong generalization to held-out rewards. For SMC-based methods, sampling particles from the posterior distribution yields significant improvements over those that sample directly from the prior, highlighting the importance of posterior-informed initialization. This improvement is particularly notable in complex tasks where high-reward outputs are rare, such as layout-to-image generation and quantity-aware generation. For example, in quantity-aware generation, negative smooth L1 loss improves from 1.804 with TDS (base SMC) to 1.077 with Top-$K$-of-NN and further to 0.850 with our $\Psi$-Sampler. Similarly, for layout-to-image generation, mIoU increases from 0.417 (TDS) to 0.425 with Top-$K$-of-NN and 0.467 with $\Psi$-Sampler. In contrast, initializing with ULA or MALA yields only marginal gains or even degraded performance, due to the lack of Metropolis-Hastings correction in ULA and the limited exploration capacity of MALA in high-dimensional spaces. Single-particle methods consistently underperform compared to SMC-based methods.
Ablation Study.
We conduct an ablation study that examines how performance varies under different allocations of a fixed total NFE between the initial particle sampling stage (via Top-$K$-of-$K$ or MCMC) and the subsequent SMC stage; full results and analysis are provided in Appendix G.
We additionally provide quantitative and qualitative results on all three applications using another score-based generative model, SANA-Sprint [29] in Appendix H.
@@ 5.3 Qualitative Results
From the first and second rows of Fig. 2, we observe that baseline methods often fail to place objects correctly within the specified bounding boxes or generate them in entirely wrong locations. For instance, in the first row, most baselines fail to position the bird accurately, and in the second row, none correctly place the car. For quantity-aware generation, the fourth row shows the counted results corresponding to the third row. While $\Psi$-Sampler successfully generates the target number of blueberries in an image, the baselines exhibit large errors—Top-$K$-of-NN comes closest but still misses some. In rows 5 and 6, only $\Psi$-Sampler correctly generates the target number of coins. In the aesthetic preference task, although all methods produce realistic images, $\Psi$-Sampler generates the most visually appealing image with the highest aesthetic score. Additional qualitative examples are provided in the Appendix I.
@@ 5.4 Evaluation of Initial Particles
In Fig.4, we compare MALA and pCNL on the layout-to-image generation task across varying step sizes using four metrics: acceptance probability, reward (mIoU via GroundingDINO [72]), held-out reward (Salience DETR [74]), and sample diversity (LPIPS MPD [79]). All metrics are directly computed from the Tweedie estimates [18] of MCMC samples, before the SMC stage. As the step size increases, MALA’s acceptance probability rapidly drops to near-zero, while pCNL maintains stable acceptance probability. Larger step sizes generally improve reward scores, with performance tapering off at excessively large steps. Held-out reward trends mirror this pattern, suggesting that the improvements stem from genuinely higher-quality samples rather than reward overfitting [60, 20]. Although LPIPS MPD slightly declines with increasing step size due to reduced acceptance, pCNL at step size 2.0 maintains diversity on par with MALA at 0.05. Additional results for other tasks are included in the Appendix F.
@@ 6 Conclusion and Limitation
7 Acknowledgements
This work was supported by the NRF of Korea (RS-2023-00209723); IITP grants (RS-2019-II190075, RS-2022-II220594, RS-2023-00227592, RS-2024-00399817, RS-2025-25441313, RS-2025-25443318, RS-2025-02653113); and the Technology Innovation Program (RS-2025-02317326), all funded by the Korean government (MSIT and MOTIE), as well as by the DRB-KAIST SketchTheFuture Research Center.
@@ Appendix A Reward Alignment with Stochastic Optimal Control
[…]
@@ Appendix C Acceptance Probability of MALA and pCNL
$\frac{\mathcal{N}(\mathbf{x};\;\boldsymbol{\mu},\mathbf{C})}{\mathcal{N}(\mathbf{x};\;\mathbf{0},\mathbf{C})}\;=\;\exp\left(-\frac{1}{2}\|\boldsymbol{\mu}\|^{2}_{\mathbf{C}}+\langle\boldsymbol{\mu},\mathbf{x}\rangle_{\mathbf{C}}\right)$
@@
$\tilde{q}_{0}(\tilde{\mathbf{x}}|\mathbf{x})$ $=\mathcal{N}(\tilde{\mathbf{x}};\;\mathbf{0},\mathbf{I})$ (35)
@@
$\frac{p_{1}^{*}(\mathbf{x}^{\prime})q_{M}(\mathbf{x}|\mathbf{x}^{\prime})}{p_{1}^{*}(\mathbf{x})q_{M}(\mathbf{x}^{\prime}|\mathbf{x})}$ $=\frac{\mathcal{N}(\mathbf{x}^{\prime};\;\mathbf{0},\mathbf{I})\exp(r(\mathbf{x}_{0|1}^{\prime})/\alpha)\mathcal{N}(\mathbf{x};\;\mathbf{x}^{\prime}(1-\epsilon/2)+\epsilon/2\cdot\nabla r(\mathbf{x}_{0|1}^{\prime})/\alpha,\epsilon\mathbf{I})}{\mathcal{N}(\mathbf{x};\;\mathbf{0},\mathbf{I})\exp(r(\mathbf{x}_{0|1})/\alpha)\mathcal{N}(\mathbf{x}^{\prime};\;\mathbf{x}(1-\epsilon/2)+\epsilon/2\cdot\nabla r(\mathbf{x}_{0|1})/\alpha,\epsilon\mathbf{I})}$ (45)
@@ Appendix D Experimental Setup and Details
We use FLUX-Schnell [28] as the score-based generative model for our method and all baselines. Although FLUX is a flow-based model, SMC-based inference-time reward alignment can be applied by reformulating the generative ODE as an SDE [7, 13], as described in Sec. 3.1. Further to ensure diversity between samples during SMC, we applied Variance Preserving (VP) interpolant conversion [13, 54]. Apart from FLUX, we additionally report quantitative and qualitative results using another score-based generative model, SANA-Sprint [29], in Appendix H. These results demonstrate that the our claims are not tied to a specific model architecture, and additionally highlighting the robustness of $\Psi$-Sampler.
We use 25 denoising steps for all SMC-based methods and 50 for single-particle methods to compensate for their reduced exploration capacity. To ensure fair comparison, we fix the total number of function evaluations (NFE) across all SMC variants—1,000 for layout-to-image and quantity-aware generation tasks, and 500 for aesthetic-preference image generation. For aesthetic-preference image generation, we used half the NFE compared to other tasks because we found it sufficient for performance convergence. For methods that sample from the posterior, half of the NFE is allocated to the initial sampling stage, resulting in 20 particles during SMC, while prior-based methods use all NFE for SMC with 40 particles. For the aesthetic-preference image generation, we use half the number of particles in both settings to reflect the halved NFE. The ablation study comparing the performance on varying NFE allocation is provided in Appendix G. For all experiments, we used NVIDIA A6000 GPUs with 48GB VRAM.
Appendix E Toy Experiment Setup and Details
In this section, we describe the setup for the toy experiment presented in the main paper. Specifically, the data distribution $p_{0}$ is a six-mode Gaussian mixture with covariance $0.3\mathbf{I}$, consisting of six equally weighted components uniformly arranged on a circle of radius 6. The reward function is defined as the sum of three Gaussian components centered at points evenly spaced on the circle: $r(x)=\sum_{i=1}^{3}\exp{(-0.5\|x-p_{i}\|^{2}})$, where $p_{1}=(6.5,0)$, $p_{1}=(-3.25,3.25\sqrt{3})$, and $p_{3}=(-3.25,-3.25\sqrt{3})$. We trained a few-step score-based generative model using 4-layer MLP with hidden dimension 128. For sampling, we fixed the total NFE to 100 across all methods and visualized the results using 2,000 generated samples. We used step size 0.1 for MALA and 0.2 for pCNL. We deliberately used a smaller step size for MALA than for pCNL to better reflect the actual experimental settings used in the main experiments (Sec. 5.1).
Appendix F Additional Evaluation Results of MALA and pCNL Initializations under Varying Step Size